OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n. The integer n3 + 2n is divisible by 3 for every positive integer n

OpenStudy (anonymous):

i take it it is \(n^3+2n\) right?

OpenStudy (anonymous):

im actually really confuded

OpenStudy (anonymous):

k, have you ever done a proof by induction?

OpenStudy (anonymous):

nope:(

OpenStudy (anonymous):

damn then we have a long way to go there is a method, and a bit of algebra will be used step one (the easy step) show that it is true if \(n=1\) i .e. replace \(n\) by \(1\) and see what you get

OpenStudy (anonymous):

what do you get if \(n=1\) ?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

yes is 3 divisible by 3?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

lol, yes! so it is true if \(n=1\) and the first step is done

OpenStudy (anonymous):

step two, assume it is true if \(n=k\) this may be confusing , but just replace \(n\) by \(k\) and let me know what you get

OpenStudy (anonymous):

wait so like k3 + 2k

OpenStudy (anonymous):

yes, \(k^3+2k\) exactly so we assume that \(k^2+2k\) is divisible by \(3\)

OpenStudy (anonymous):

with that assumption, (now comes the only hard part) we want to show that it is true if \(n=k+1\) replace \(n\) by \(k+1\)and let me know what you get

OpenStudy (anonymous):

k+1=k+1?

OpenStudy (anonymous):

replace \(n\) in the expression \(n^3+2n\) by \(k+1\)

OpenStudy (anonymous):

okay k+1^3+2(k+1)

OpenStudy (anonymous):

right now comes the algebra

OpenStudy (anonymous):

can you expand this nonsense \[(k+1)^3+2(k+1)\]?

OpenStudy (anonymous):

(k+1)(k+1)(k+1)+2k+2?

OpenStudy (anonymous):

yeah, that is step one

OpenStudy (anonymous):

i don't mean to torture you \[(k+1)^2=k^3+3k^2+3k+1\]

OpenStudy (anonymous):

so you have \[k^3+3k^2+3k+1+2k+2\]

OpenStudy (anonymous):

oh no scratch that lets take a \(k^3+2k\) out of this and write it as \[k^3+2k+3k^2+3k+3\]

OpenStudy (anonymous):

or better yet \[\color{red}{k^3+2k}+\color{blue}{3(k^2+k+1)}\]

OpenStudy (anonymous):

we know that \(\color{red}{k^3+2k}\) is divisible by 3 "by induction" i.e. it is what we assumed to be true at the beginning of step 2 and we know that \(\color{blue}{3(k^2+k+1)}\) is divisible by 3 because of that 3 out front therefore the whole thing has to be divisible by 3, and that is what you had to prove

OpenStudy (anonymous):

WOW THANK YOU SO MUCH ♥